OpenStudy (anonymous):

Hi guys. I am looking for the equation of the tangent line to y=4sin(x) at x=3pi/4. I took the deriv of 4sin(x) and got 4cos(x). plugged it into the equation and got: y-(4(sqrt2)/2 = (-4(sqrt2)/2)(x-(3pi/4). now I am stuck.......suggestions?

OpenStudy (anonymous):

\[y-(4\sqrt{2)}/2 = -((4\sqrt{2)}/2)(x-3\pi/4)\]